OpenStudy (thefluffmuncher):

Solve: 7/9=21/x A. 14 B. 27 C. 14/9 D. 3

OpenStudy (mathlegend):

\[\frac{ 7 }{ 9 } * \frac{ a }{ b} = \frac{ 21 }{ x }\] What do you multiply 7 times to get 21?

OpenStudy (thefluffmuncher):

3

OpenStudy (mathlegend):

What you do to the top (numerator) you have to do to the bottom (denominator)... right? So multiply the bottom by 3 as well.

OpenStudy (thefluffmuncher):

thanks dude :D you are really a math legend

OpenStudy (mathlegend):

I try :)